use the following f(x)=2x^2-3, g(x)=2x-3 for the following: (fg)(3)
Answer by nerdybill(7384)

use the following f(x)=2x^2-3, g(x)=2x-3 for the following: (fg)(3)
.
(fg)(x) = 2(2x-3)^2-3
(fg)(3) = 2(2(3)-3)^2-3
(fg)(3) = 2(6-3)^2-3
(fg)(3) = 2(3)^2-3
(fg)(3) = 2(9)-3
(fg)(3) = 18-3
(fg)(3) = 15